OpenStudy (anonymous):

Find the inverse of the function. f(x) = the cube root (x /7 )- 9

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{\frac{ x }{ 7 } - 9}\] Is This The Equation?

OpenStudy (anonymous):

No the cube root is only for x/7

OpenStudy (anonymous):

thecollinjc1 you still there

OpenStudy (anonymous):

So Are You Looking For Something Like This ?

OpenStudy (anonymous):

Kind of here are the possible answers Find the inverse of the function. f-1(x) = 21(x + 9) f-1(x) = [7(x + 9)]^3 f-1(x) = 7(x^3 + 9) f-1(x) = 7(x + 9)^3

OpenStudy (anonymous):

Not Positive on This One But I'd Guess The Last One. @jim_thompson5910 Help?

OpenStudy (anonymous):

I have one more question after this one if we can solve this one if want to help me on it.

OpenStudy (jdoe0001):

\(\bf \Large f(x) = y=\sqrt[3]{\cfrac{ x }{ 7 } }- 9\\ f^{-1}(x) = y^{-1} \implies \Large x=\sqrt[3]{\cfrac{ y }{ 7 } }- 9\)

OpenStudy (jdoe0001):

you just switch the variables, and then solve for "y"

OpenStudy (anonymous):

What do you mean switch the variables?

OpenStudy (jdoe0001):

\(\bf \bf \Large f(x) = \color{red}{y}=\sqrt[3]{\cfrac{ \color{blue}{x} }{ 7 } }- 9\\ f^{-1}(x) = y^{-1} \implies \Large \color{blue}{x}=\sqrt[3]{\cfrac{ \color{red}{y} }{ 7 } }- 9\)

OpenStudy (anonymous):

DO I multiply them I will fan you if you tell me what the answer is by explaining it step by step